# DECO: Life-Cycle Management of Enterprise-Grade Chatbots

### v1

@@ 10. Conclusion
In this paper, we introduced DECO, a comprehensive framework designed for the end-to-end lifecycle management of enterprise-grade chatbots tailored to software engineering tasks. DECO integrates robust preprocessing pipelines with seamlessly connected front-end and back-end modules, utilizing mature Azure solutions to facilitate full-scale deployment and monitoring in production environments. The system utilizes data from historically resolved incident tickets to bridge knowledge gaps, thereby creating a rich documentation dataset rooted in prior engineering practices. Its hierarchical skill selection design ensures that the most relevant sources of context and internal tools from various teams are employed, utilizing advanced retrieval algorithms to efficiently identify pertinent incidents and documents. A continuous evaluation pipeline enables DECO to refine its retrieval algorithms and response generation constantly, guided by real-world usage and feedback. Deployed across Microsoft, DECO has demonstrated its effectiveness by reducing the time engineers spend on incident management by 10-20 minutes per incident, significantly enhancing engineering workflows. Future developments for DECO will focus on further improving the algorithm, enhancing document retrieval accuracy based on user feedback, and improve memory management techniques to further reduce the context length with longer chat history or infer user preferences.

### v2

@@ 10. Conclusion
In this paper, we introduced DECO, a framework for end-to-end management and deployment of copilots tailored to software engineering tasks. DECO integrates preprocessing pipelines with tightly connected front-end and back-end modules to support multi-turn interactions with users. Leveraging historical incident tickets, DECO bridges knowledge gaps by enriching documentation with insights from past engineering practices. Its hierarchical skill selection with agentic design dynamically selects the most relevant context and tools, using advanced retrieval algorithms to surface pertinent incidents, documents, and code snippets. A continuous evaluation pipeline monitors retrieval performance and response quality. Deployed across Microsoft, DECO now serves over 500 Monthly Active Users (MAUs), reducing time spent on incident management and improving engineering workflows. Future work will focus on refining retrieval algorithms, improving document relevance through user feedback, and enhancing memory management to handle longer chat histories more efficiently.